Suppose that 30% of all college students smoke cigarettes. A sample of 16 is selected randomly.

. What is the probability that between 5 and 8 (both inclusive) students smoke? Round your answer to four decimal places.
Answer #1

p = Proportion of college students that smoke cigarettes = 0.3

n = Sample size = 16

Let X denote the number of students in the sample that smoke cigarettes.

X \sim Binomial(n=16,p=0.3)

Probability that between 5 and 8 students smoke

= P[5 \leq X \leq 8]

= P(X=5) + P(X=6) + P(X=7) + P(X=8)

= \binom{16}{5}0.3^5 0.7^{11} + \binom{16}{6}0.3^6 0.7^{10} + \binom{16}{7}0.3^7 0.7^{9} + \binom{16}{8}0.3^8 0.7^{8}= 0.2099 + 0.1649 + 0.1010 + 0.0487

= 0.5245 (Answer)